A nitrogen gas container has a volume of 35.0 L. Calculate the mass of nitrogen in the container if the gas has a pressure of 1800 torr at 42°C.

1
Expert's answer
2021-11-11T08:07:01-0500

"V=35.0 \;L \\

p=1800 \;torr = 2.368 \;atm \\

T= 42 + 273 = 315 \;K"

Ideal Gas Law

pV=nRT

R = 0.08206 L×atm/mol×K

"n= \frac{pV}{RT} \\

n = \frac{2.368 \times 35.0}{0.08206 \times 315} = 3.206 \;mol"

M(N2) = 28.01 g/mol

"m = 3.206 \times 28.01 = 89.81 \;g"

Answer: 89.81 g